### v1

@@ 4.3 Model Validation
We assess the model’s empirical validity by testing its ability to predict both the level and heterogeneity of price and quantity changes following the 2017–2018 tariff increases. These moments were not targeted in the estimation.
Model-predicted price changes are computed as:
$\widehat{\Delta\ln p_{ijht}}=\Phi_{ijht}(s_{ijht},x_{ijht}\mid\boldsymbol{\hat{\Theta}})\cdot\Delta\ln(1+\tau_{cht}),$ (4.4)
where $\tau_{cht}$ denotes the ad-valorem tariff on product $h$ from country $c$, and $\Phi_{ijht}$ is the model-implied pass-through elasticity defined in equation (2.11), which depends on bilateral shares and the estimated parameter vector $\boldsymbol{\hat{\Theta}}$.
Although the model is primarily designed to explain prices, it embeds a demand-driven allocation rule that links prices to quantities (and sales) via the importer’s demand curve. This mapping imposes a specific quantity-setting structure, enabling a direct comparison between predicted price responses and observed adjustments in quantities and trade values. Predicted quantity changes are given by $\widehat{\Delta\ln q_{ijht}}=-\hat{\varepsilon}_{ijht}\cdot\widehat{\Delta\ln p_{ijht}},$ where $\hat{\varepsilon}_{ijht}$ denotes the match-specific residual demand elasticity implied by the model. For sales, the corresponding mapping is $\widehat{\Delta\ln r_{ijht}}=-(1-\hat{\varepsilon}_{ijht})\cdot\widehat{\Delta\ln p_{ijht}}$.
@@
Panel A: Data
(1) (2) (3) (4)
$\Delta\ln(1+\tau_{cht})$0.151 -0.066 -0.223 -0.163
(0.093) (0.093) (0.109) (0.107)
$\Delta\ln(1+\tau_{cht})\cdot s_{ijht-1}$ 0.054 0.029
(0.072) (0.157)
$\Delta\ln(1+\tau_{cht})\cdot x_{ijht-1}$0.403 -0.271
(0.113) (0.135)
R-squared 0.04 0.04 0.31 0.31
Panel B: Model
(1) (2) (3) (4)
$\Delta\ln(1+\tau_{cht})$0.248 -0.144 -0.249 -0.136
(0.008) (0.009) (0.010) (0.011)
$\Delta\ln(1+\tau_{cht})\cdot s_{ijht-1}$ 0.104 0.091
(0.011) (0.010)
$\Delta\ln(1+\tau_{cht})\cdot x_{ijht-1}$0.475 -0.486
(0.026) (0.029)
R-squared 0.32 0.46 0.50 0.59
$\text{FE}_{ht}+\text{FE}_{cs}$ Yes Yes No No
$\text{FE}_{ht}+\text{FE}_{ct}+\text{FE}_{jt}$ No No Yes Yes
@@
Notes: This table reports the pass-through of tariffs to duty-exclusive prices at the exporter–importer–product level. Panel A presents reduced-form estimates from the data. Panel B shows corresponding pass-through estimates generated by the model. Columns (2)–(4) interact tariff changes with lagged supplier share ($s_{ijht-1}$) and lagged buyer share ($x_{ijht-1}$). Columns (1) and (2) use baseline fixed effects ($\text{FE}_{ht}+\text{FE}_{cs}$), while Columns (3) and (4) employ a more stringent specification with product–year, country–year, and buyer–year fixed effects ($\text{FE}_{ht}+\text{FE}_{ct}+\text{FE}_{jt}$). Standard errors are clustered at the HS8 product and exporter-country level. Observation counts are rounded to four significant digits per U.S. Census Bureau disclosure guidelines. Source: FSRDC Project Number 2109 (CBDRB-FY25-P2109-R12520).
4.3.1 Price Predictions
We begin with price outcomes. Table 5 compares tariff pass-through elasticities in the data (Panel A) and in the model (Panel B), in terms of average effects and their heterogeneity with respect to buyer and supplier shares. Columns (1) and (2) adopt a baseline specification with product–time and country–sector fixed effects, while Columns (3) and (4) adopt the more demanding specification with buyer–time and country–time fixed effects.
The model predicts an average pass-through elasticity on duty-exclusive prices of $-0.25$, which translates to a 75% pass-through rate (Columns (1) and (3) of Panel B). These values fall within the 95% confidence intervals of the corresponding reduced-form estimates in the data (Columns (1) and (3) of Panel A).
Columns (2) and (4) of Panel A, consistent with Column (6) of Table 3, show that pass-through declines significantly with buyer share ($x_{ijht}$) but exhibits no robust relationship with supplier share ($s_{ijht}$). The model closely replicates these patterns: as shown in Panel B, pass-through declines steeply in $x_{ijht}$ and increases mildly in $s_{ijht}$, with magnitudes comparable to those in the data. This alignment supports our interpretation of Table 3 as consistent with a low-$\theta$, high-$\phi$ environment, which is also consistent with our structural estimates.
4.3.2 Goodness-of-fit Test
Having shown that the model replicates average and heterogeneous price responses to tariff shocks, we formally test its predictive performance by evaluating how well model-implied price changes explain observed variation. This exercise complements Table 5 by providing a direct measure of goodness of fit relative to standard alternatives.
Specifically, we estimate:
$\Delta\ln p_{ijht}=\beta\,\,\widehat{\Delta\ln p_{ijht}}\,+\,\mathbf{FE}\,+\,u_{ijht},$ (4.5)
where $\Delta\ln p_{ijht}$ is the observed change in the duty-inclusive price for product $h$ between exporter $i$ and importer $j$, and $\widehat{\Delta\ln p_{ijht}}$ is the corresponding model-predicted change.
To benchmark the results, we compare the full model to three nested alternatives that sequentially shut down bilateral bargaining ($\phi=0$), decreasing returns to scale ($\theta=1$), or both. The fully restricted case ($\phi=0$, $\theta=1$) corresponds to a standard Nash-Bertrand model with constant marginal costs (e.g., Dhyne et al., 2022); the intermediate case with $\phi>0$ and $\theta=1$ mirrors the bargaining framework of Gopinath and Itskhoki (2011), which overlooks the heterogeneous buyer dimension. In all cases, we use the same parameter values from Column (1) of Table 4, without re-estimating. [^31] As discussed in Section 4.1, $\theta$ is not identified when $\phi=0$. Moreover, fixing $\theta=1$ and estimating equation (4.5) using the value of $\phi$ jointly obtained with $\theta$ from equation (4.2) yields an upper bound on the value of $\hat{\beta}$ attainable with a re-estimated $\phi$. Appendix E.3 provides formal proof and supporting simulation evidence.
A key challenge is that observed price changes may reflect shocks unrelated to tariffs, increasing the risk of rejecting a valid model for reasons unrelated to its tariff predictions. In addition, while the model uses statutory tariff changes, observed duty-inclusive prices reflect actual duties paid, potentially introducing measurement error. To address both issues, we follow the logic of the IV-based goodness of fit test in Adão et al. (2023) and estimate equation (4.5) using two-stage least squares, instrumenting $\widehat{\Delta\ln p_{ijht}}$ with statutory tariff changes. This isolates variation in predicted prices that is directly attributable to tariff shocks, which is the one targeted by the model. Under the null that the model accurately captures pass-through, the IV coefficient $\hat{\beta}$ should equal one.
Figure 3: IV-Based Goodness-of-Fit Test
Notes: Each point reports the coefficient from an IV regression of observed log price changes on model-predicted changes $\widehat{\Delta\ln p_{ijht}}$, using statutory tariffs as instruments. Lines show 95% confidence intervals. Blue and red denote regressions with $\mathrm{FE}_{ht}+\mathrm{FE}_{cs}$ and $\mathrm{FE}_{ht}+\mathrm{FE}_{ct}+\mathrm{FE}_{jt}$, respectively. Standard errors are clustered by product and exporter-country. Observation counts (249,000) are rounded per Census disclosure guidelines. Source: FSRDC Project Number 2109 (CBDRB-FY25-P2109-R12520).
Figure 3 presents the results across the four model variants, each estimated under two alternative fixed effects specifications. Blue points and lines correspond to regressions with product–time and country–sector fixed effects, while red ones use the more stringent specification with product–time, country–time, and buyer–time fixed effects. In all cases, the baseline model provides the best fit: the estimated coefficient is very close to one and not statistically different from it under both specifications.
Models with constant returns to scale ($\theta=1$), whether or not they include bargaining, perform noticeably worse, although their coefficients are not rejected at conventional significance levels. These variants lack a cost channel, and while the case with $\phi=0$ features strategic complementarities that generate some degree of incomplete pass-through, this mechanism alone does not produce sufficient variation to match the data. Similarly, the fourth model with decreasing returns but no bargaining ($\phi=0$, $\theta<1$), which was previously rejected based on price-level evidence (Table 2), also underperforms relative to the baseline model. In this case, markup responses driven by strategic complementarities amplify cost-based adjustments, but in a way that also fails to replicate the observed pass-through patterns. [^32] Appendix D.4 (Table D.6) presents additional robustness checks using alternative calibrations, including $\rho=5$ instead of 10 and $\varrho=0.5$ instead of 1. The model’s predictive performance remains stable across these variations, lending further support to our main findings.
Overall, these results suggest that cost adjustments, along with weak strategic complementarities, are necessary to account for the empirical evidence. This is consistent with a setting in which oligopsony forces are dominant.
4.3.3 Quantity Predictions
Next, we assess the model’s ability to predict changes in bilateral quantities. Table D.7 in Appendix D.4 reports relationship-level quantity responses to tariff changes. Columns (3)–(4) of Panel A show that tariff increases reduce traded volumes, with larger declines under more stringent fixed effects. However, the interaction terms with supplier and buyer shares are statistically imprecise, suggesting inconclusive evidence of heterogeneity in the data. [^33] In robustness exercises (not shown), the signs on the interaction coefficients with supplier and buyer shares occasionally flip, though they largely remain statistically insignificant. We therefore conclude that the data do not offer robust evidence of heterogeneous quantity responses by relationship structure.
Panel B shows that the model generates sizable average quantity declines and predicts heterogeneity across relationships. In particular, quantity responses become less negative with higher buyer shares, consistent with the model’s allocative logic: stronger buyers face smaller price increases and thus smaller quantity reductions. In contrast, the positive interaction with supplier share is not explained by price adjustments because the model predicts little variation in pass-through along this dimension. Instead, it reflects the curvature of the importer’s demand curve embedded in the model.
Table D.8 in Appendix D.4 evaluates model fit for quantities (Panel A) and sales (Panel B) across the four alternative parameterizations discussed above. Since all models share the same demand system, performance differences reflect variation in price predictions rather than differences in quantity mechanisms. While all specifications are formally rejected, the baseline model and the specification with decreasing returns to scale perform best, mirroring their superior performance in prices.
Despite differences in magnitude and precision, the model captures the broad directional patterns in quantity responses, lending support to its core allocative mechanism. However, its weaker quantitative fit suggests that additional forces beyond price-based allocation shape the observed quantity adjustments. The model’s strong performance on prices highlights this asymmetry and points to potential gains from extending the framework to incorporate supply-side responses.
5 Aggregate Implications of Bargaining in Firm-to-Firm Trade
This section examines the impact of the Trump tariffs on aggregate import prices through the lens of our bargaining model. We begin by situating our analysis within the broader literature on tariff pass-through. We then use the model to simulate the aggregate effects of the tariffs and assess the mechanisms driving price adjustments.
5.1 Comparison with Existing Pass-Through Estimates
Several recent studies find near-complete pass-through of the 2018 Trump tariffs to U.S. import prices (e.g., Fajgelbaum et al., 2020; Amiti et al., 2019b, 2020). These analyses primarily rely on product-level data. In contrast, our match-level estimates indicate incomplete pass-through. For instance, in a standard specification similar to those used in the literature, Column (1) of Table 5 (Panel A) reports a pass-through elasticity on duty-exclusive prices of $-0.15$, implying an 85% pass-through rate.
This divergence partly reflects differences in sample composition. As discussed in Section 3.3, our estimates are based on repeated firm-to-firm relationships where bilateral prices can be observed over time. They therefore capture within-relationship price changes net of compositional shifts across firms or products, rather than on changes in product-level unit values. In addition, we restrict attention to arm’s-length transactions involving intermediate goods and suppliers with two or more buyers. Within this sample, we estimate pass-through rates ranging from 78 to 85%, depending on the fixed effects used, as shown in Table 5. The model’s predicted pass-through closely matches these empirical patterns.
To further assess the role of sample composition, Table D.9 in Appendix D.5 reports pass-through estimates across alternative sample definitions. Expanding the baseline to include single-buyer relationships raises the estimate to approximately 83%. Adding matches involving related parties, energy goods, or extreme price changes increases it to roughly 90%. The most inclusive specification, which further incorporates final consumption goods, yields pass-through elasticities between 93% and 95%, depending on the fixed effects. Although these broader samples imply higher pass-through, they still fall short of full pass-through, in contrast to product-level studies. [^34] Using the monthly LFTTD data, we show in Table D.10 in Appendix D.5 that pass-through remains incomplete in product-level regressions when restricting to the firms and products used in our baseline sample, confirming that this is a feature of repeated firm-to-firm relationships.
These product-level estimates are based on aggregated unit values that combine within-match price changes with shifts in the composition of transactions, particularly the inclusion of one-off (spot) exchanges. Such transactions may differ systematically from repeated firm-to-firm relationships. In particular, incomplete pass-through may be a feature of relational trade, where prices reflect bilateral bargaining and upward-sloping residual supply. Spot exchanges, by contrast, are more likely to reflect marginal cost pricing and thus exhibit full pass-through. While speculative, this interpretation is consistent with Cajal-Grossi et al. (2023), who find lower markups and more competitive pricing in spot relationships.
5.2 Tariffs and Aggregate Import Prices
We use the estimated model to assess the impact of the 2018 tariffs on aggregate import prices and decompose the underlying contributions of markup and marginal cost changes. A full evaluation of the tariffs’ macroeconomic effects would require a general equilibrium framework that incorporates changes in expenditure, wages, export prices, and retaliation, as in Fajgelbaum et al. (2020). While such elements are essential for welfare analysis, they lie beyond the scope of our model.
Instead, we focus on one central component of tariff incidence: the change in aggregate import prices within ongoing firm-to-firm relationships. We compute model-implied bilateral price changes using equation (4.4) and estimate aggregate pass-through by regressing these predicted changes on the tariff shock, weighting each observation by its initial-period import value. The approach connects the model’s micro-level predictions to aggregate outcomes and provides a framework for decomposing the contribution of distinct adjustment channels.
To that end, we separately isolate the roles of markup and marginal cost adjustments. For each mechanism, we compute the model-implied price changes while holding the other channel constant, and re-estimate the weighted regression. The predicted price change due solely to markup adjustment is defined as:
$\widehat{\Delta^{\Gamma}\ln p_{ijht}}=\Phi^{\Gamma}_{ijht}(s_{ijht},\,x_{ijht}\,\mid\,\hat{\boldsymbol{\Theta}})\cdot\Delta\ln T_{cht},\quad\text{where}\quad\Phi^{\Gamma}_{ijht}\equiv\frac{1}{1+\Gamma_{ij}}.$
Similarly, the predicted price change driven solely by cost adjustment is given by:
$\widehat{\Delta^{\Lambda}\ln p_{ijht}}=\Phi^{\Lambda}_{ijht}(s_{ijht},\,x_{ijht}\,\mid\,\hat{\boldsymbol{\Theta}})\cdot\Delta\ln T_{cht},\quad\text{where}\quad\Phi^{\Lambda}_{ijht}\equiv\frac{1}{1+\Lambda_{ij}}.$
@@ 5.2 Tariffs and Aggregate Import Prices
Panel A: Aggregate Passthrough ()
Agg. pass-through elasticity 70.7 64.6
Cost channel only: $1/(1+\Lambda_{ij})$ 69.3 62.7
Markup channel only: $1/(1+\Gamma_{ij})$ 97.1 88.0
@@
Cost Elasticity: $\Lambda_{ij}$ 1.01 0.99
Markup Elasticity: $\Gamma_{ij}$0.01 0.01
@@
Notes: This table reports model-implied aggregate pass-through estimates following the 2018 U.S. tariff increases. Column (1) includes product–time and country–sector fixed effects. Column (2) includes product–time, country–time, and buyer–time fixed effects. The overall pass-through elasticity is computed as $1+$ the estimated coefficient on $\Delta\ln(1+\tau)$, and decomposed into contributions from the cost channel ($\Lambda_{ij}$) and the markup channel ($\Gamma_{ij}$). The counterfactual “Cost channel only” row shows the predicted pass-through when markup elasticities are set to zero, while the “Markup channel only” row sets cost elasticities to zero. Panel B reports the relative contribution of each channel to the cross-sectional variance of $\Lambda_{ij}+\Gamma_{ij}$, the total elasticity governing pass-through. These shares sum to one and are derived from a variance decomposition. Source: FSRDC Project Number 2109 (CBDRB-FY25-P2109-R12520).
@@
Panel A of Table 6 reports the aggregate tariff pass-through predicted by the model, which incorporates both markup and marginal cost adjustments: 71% under the baseline specification (Column (1)) and 65% under the alternative using product–time, country–time, and buyer–time fixed effects (Column (2)). These estimates are lower than the average pass-through of 75% obtained from unweighted regressions (Table 5), reflecting the more incomplete pass-through associated with larger, high-share buyers who account for a greater share of aggregate trade.
Nearly all of the predicted price response is driven by the cost channel. When markup elasticities are shut down (i.e., setting $\Gamma_{ij}=0$), the model yields pass-through rates between 69 and 63%, closely matching the full-model estimates. By contrast, shutting down the cost elasticities results in much higher pass-through, ranging from 88 to 97%.
Panel B offers an alternative approach to quantify the contribution of each channel by decomposing the variance of the term $\Lambda_{ij}+\Gamma_{ij}$, which governs the pass-through elasticity $1/(1+\Lambda_{ij}+\Gamma_{ij})$. This decomposition attributes the cross-sectional variation in pass-through to cost and markup elasticities, with their shares summing to one by construction. The results reinforce the earlier findings: nearly all of the variation is explained by the cost elasticity $\Lambda_{ij}$, underscoring its dominant role in shaping price responses.
Together, the results suggest a clear conclusion: tariff pass-through is incomplete, reflecting the combined effects of cost-side adjustments and bargaining. The underlying mechanism is intuitive. When tariffs reduce demand from large U.S. buyers, exporters move down their marginal cost curves, dampening the price impact of the shock. Although markups adjust endogenously, the opposing forces of strategic complementarities and substitutabilities limit their quantitative contribution to aggregate price change. This is a robust implication of our framework. As illustrated in Figure 2, when returns to scale are decreasing ($\theta<1$) and buyer power is high (large $\phi$), most of the variation in pass-through across the $(s_{ij},\,x_{ij})$ space is driven by cost adjustments, with a comparatively smaller role for markup responses.
5.3 The Importance of Bargaining in Firm-to-Firm Trade
Although bargaining plays a central role in our model, it appears to have limited influence on tariff pass-through, which primarily reflects cost-side adjustments. This is not because bargaining is unimportant, but because it endogenously mutes the markup channel. When buyer power is strong ($\phi$ high), the markup elasticity $(\Gamma_{ij})$ tends toward zero. As a result, pass-through is driven almost entirely by cost adjustments and appears relatively unresponsive to markup variation. In this sense, the weak role of the markup channel in shaping pass-through is itself a consequence of strong importer market power.
In addition, bargaining remains essential to understand price levels, markups, and the welfare consequences of market power. As shown in Table 2, two-sided market power is critical to explaining the cross-sectional variation in prices and markups. Moreover, as discussed in Section 2.5, $\phi$ governs how markup dispersion maps into misallocation. When exporters hold bargaining power, markup heterogeneity leads to underproduction. But when bargaining power shifts toward importers, the same dispersion can improve efficiency by reallocating output toward lower-cost suppliers. Ignoring bargaining would therefore yield misleading conclusions about the allocative effects of firm-to-firm trade.

### v2

@@ 4.3 Model Validation
Proposition 2 yields a closed-form expression for the pass-through elasticity, which we use to assess the model’s out-of-sample performance. Specifically, we examine whether the model can predict the level and heterogeneity of price responses to the 2017–2018 tariff increases, which are excluded from estimation. Using observed shares in 2017 and the estimated parameters, the model-implied change in duty-exclusive prices is
$\widehat{\Delta\ln p^{*}_{ijht}}=\Phi^{*}_{ijht}\!\left(s_{ijht},x_{ijht}\mid\boldsymbol{\hat{\Theta}}\right)\cdot\Delta\ln\left(1+\tau_{c(i)ht}\right),$ (4.5)
where $\tau_{cht}$ denotes the statutory tariff rate and $\Phi^{*}_{ijht}$ is the model-implied pass-through elasticity into duty-exclusive prices. [^29] Proposition 2 characterizes pass-through into duty-inclusive prices; pass-through into duty-exclusive prices follows as $\Phi^{*}=\Phi-1$. Because the model delivers a partial-equilibrium mapping from tariffs to prices while the data reflect equilibrium outcomes, this exercise constitutes a joint test of the model’s partial-equilibrium structure and its underlying mechanisms.
Table 6 reports pass-through elasticities in the data and in the model for the baseline specification (equation (3.2)) and three restricted variants, using baseline fixed effects. [^30] For the restricted variants, equation (4.5) is evaluated using the baseline parameter estimates from column (1) of Table 4 without re-estimation, except where parameters are restricted by construction (e.g., $\phi=0$ or $\theta=1$). Column (1) reports the empirical estimates. Column (2) corresponds to the baseline model. Column (3) shuts down both bargaining and decreasing returns by imposing $\phi=0$ and $\theta=1$, yielding a standard Nash–Bertrand model with constant marginal costs, as in kikkawa2019imperfect. Column (4) sets $\theta=1$, retaining bargaining but eliminating decreasing returns, as in gopinath2011search. Column (5) sets $\phi=0$, eliminating importer bargaining power by assigning all pricing power to exporters, while preserving decreasing returns.
@@
Panel A: Main Effect
Data Baseline $\phi=0$, $\theta=1$ $\theta=1$ $\phi=0$
(1) (2) (3) (4) (5)
$\Delta\ln(1+\tau_{cht})$ -0.126 -0.222 -0.094 -0.013 -0.279
(0.091) (0.008) (0.004) (0.001) (0.008)
R-squared 0.04 0.31 0.22 0.21 0.34
Panel B: Including Interactions
Data Baseline $\phi=0$, $\theta=1$ $\theta=1$ $\phi=0$
(1) (2) (3) (4) (5)
$\Delta\ln(1+\tau_{cht})$ -0.032 -0.145 -0.049 -0.005 -0.219
(0.091) (0.008) (0.009) (0.001) (0.009)
$\Delta\ln(1+\tau_{cht})\cdot s_{ijht-1}$ 0.033 0.125 -0.079 -0.011 0.122
(0.080) (0.013) (0.014) (0.003) (0.011)
$\Delta\ln(1+\tau_{cht})\cdot x_{ijht-1}$ -0.407 -0.421 -0.046 -0.013 -0.366
(0.112) (0.023) (0.023) (0.003) (0.022)
R-squared 0.04 0.44 0.23 0.23 0.41
$FE_{ht}+FE_{cs}$ Yes Yes Yes Yes Yes
@@
Notes: This table reports the response of duty-exclusive CIF prices to tariff changes at the exporter–importer–product level. Column (1) presents reduced-form estimates using observed price changes in the data. Columns (2)–(5) report the corresponding responses implied by model-predicted price changes under each specification. Panel A reports the main effect of tariff changes, while Panel B allows the response to vary with lagged supplier shares ($s_{ijht-1}$) and buyer shares ($x_{ijht-1}$). All specifications include product–time and country–sector fixed effects ($FE_{ht}+FE_{cs}$). Standard errors are clustered at the HS8 product and exporter-country level. Observation counts are rounded to four significant digits per U.S. Census Bureau disclosure guidelines. Source: FSRDC Project Number 2109 (CBDRB-FY25-P2109-R12520).
Panel A of Table 6 reports average pass-through of about 87% in the data. The baseline model predicts a lower pass-through, around 78%, yet remains close to the data with overlapping confidence intervals. The restricted variants also generate incomplete pass-through, to varying degrees: shutting down both bargaining and decreasing returns ($\phi=0$, $\theta=1$) yields about 91%; eliminating importer bargaining alone ($\phi=0$) produces the lowest pass-through, around 72%; removing decreasing returns alone ($\theta=1$) raises pass-through substantially. While all variants generate incomplete pass-through, confidence intervals do not reject any of them, limiting the informativeness of these differences for model fit.
As illustrated in Figure 2, pass-through reflects the balance of the cost and markup channels, and different $(\phi,\theta)$ combinations can generate similar averages through different mechanisms, particularly in the relevant range of buyer and supplier shares. In the baseline, strategic substitutabilities partly offset the cost channel (Column (2)); setting $\phi=0$ shifts the markup channel toward complementarities, further lowering pass-through and moving the model farther from the data (Column (5)); eliminating the cost channel alone still yields incomplete pass-through under strategic complementarities (Column (3)); and removing both the cost channel and the strategic complementarities produces near-complete pass-through (Column (4)). Average pass-through is therefore a weak discriminator across models, with more informative evidence coming from cross-sectional moments.
Panel B of Table 6 interacts tariff changes with lagged supplier and buyer shares. In the data, pass-through declines sharply with buyer share and increases modestly with supplier share. The baseline model closely replicates both gradients, with magnitudes comparable to those in the data. This alignment also sheds new light on Figure D.1: the approximately monotonic decline in pass-through across buyer-share quartiles, which we documented but did not interpret structurally in Section 3.4, finds a natural counterpart in the model, where the cost channel scales linearly with $x_{ij}$ (Equation 2.16).
Eliminating decreasing returns ($\theta=1$, Columns (3)–(4)) substantially attenuates the buyer-share gradient and counterfactually predicts lower pass-through for exporters with larger supplier shares. By contrast, setting $\phi=0$ (Column (5)) yields cross-sectional patterns similar to the baseline. This is expected, given Figure 2: when decreasing returns are strong, pass-through heterogeneity is largely governed by the cost channel, leaving limited scope for bargaining power to shape the cross-sectional gradient.
Taken together, the baseline is the only specification consistent with both panels. Column (3) comes closest to the data on average (91%), but is strongly rejected on cross-sectional grounds. Column (5) matches the cross-sectional patterns but noticeably overstates price absorption relative to both the data and the baseline. The model’s ability to reproduce observed patterns of tariff pass-through lends support to its underlying mechanisms and suggests that the partial equilibrium framework captures key forces governing price adjustment in firm-to-firm trade.
Additional Validation.
Table 6 uses baseline fixed effects throughout; Table D.7 in Appendix D.4 repeats the exercise under more stringent fixed effects and yields the same conclusions, both in terms of average pass-through and cross-sectional patterns.
One concern is that observed price changes may reflect shocks unrelated to tariffs, potentially leading the data to reject (or accept) a correctly (or incorrectly) specified model for spurious reasons. To address this, we formalize the comparison using the IV goodness-of-fit test of adao2023putting, which isolates tariff-driven variation in observed prices. Figure D.2 in Appendix D.4 reports the results. Consistent with Panel A of Table 6, none of the variants can be formally rejected based on average pass-through alone. Among the non-rejected specifications, the baseline provides a closer fit than Column (5), consistent with Table 6. Table D.8 shows that these results are robust to alternative calibrations of the model parameters and remain stable across various fixed-effects specifications.
Finally, we assess the model’s ability to predict bilateral quantity and sales responses. [^31] Predicted quantity and sales changes are given by $\widehat{\Delta\ln q_{ijht}}=-\hat{\varepsilon}_{ijht}\widehat{\Delta\ln p_{ijht}}$ and $\widehat{\Delta\ln r_{ijht}}=-(1-\hat{\varepsilon}_{ijht})\widehat{\Delta\ln p_{ijht}}$, where $\hat{\varepsilon}_{ijht}$ is the match-specific residual demand elasticity computed from observed shares and estimated parameters. Table D.9 in Appendix D.4 reports the results for quantities. Across specifications, tariff increases are associated with declines in traded volumes in both the data and the model, with larger effects under more demanding fixed effects. Interaction coefficients with buyer and supplier shares are imprecisely estimated in the data and sensitive to specification, suggesting limited power to detect heterogeneous effects. By contrast, the model predicts systematic patterns. Specifically, quantity responses become less negative at higher buyer shares, reflecting the smaller price increases faced by high-share buyers.
Appendix Table D.10 extends the IV tests to quantities and sales jointly. Consistent with our price results, the baseline and decreasing-returns specifications provide the best fit to the data, although these tests formally reject all model variants. This comparatively weaker fit for quantities does not necessarily undermine the model’s mechanism. Quantities are measured with significantly greater error in customs data, and the mapping from prices to quantities relies on estimated demand elasticities that may amplify this noise. Despite these measurement challenges, the results remain qualitatively consistent with the model’s core implication: prices govern the allocation of trade, with quantities adjusting along the demand curve.
5 Aggregate Pass-Through in Firm-to-Firm Trade
We use the estimated model to quantify the price impact of the 2018 tariffs within continuing buyer-supplier pairs and its underlying mechanisms. To aggregate the model’s micro-level predictions, we compute model-implied bilateral price changes and estimate a weighted pass-through for this subset of trade by regressing predicted price changes on the tariff shock, weighting observations by their initial import shares. We then use the model’s expression for cost and markup elasticities to decompose predicted price changes into its two underlying determinants. For each channel, we recompute counterfactual price changes while holding the other component fixed, and re-estimate the weighted regression.
@@ 5 Aggregate Pass-Through in Firm-to-Firm Trade
Panel A: Aggregate Pass-through (% )
Agg. pass-through: $1/(1+\Lambda_{ij}+\Gamma_{ij})$ 72.7 66.5
  Cost channel only: $1/(1+\Lambda_{ij})$ 71.5 64.8
@@
Cost Elasticity: $\Lambda_{ij}$ 1.023 0.997
Markup Elasticity: $\Gamma_{ij}$ -0.023 0.003
@@
Notes: This table reports model-implied aggregate pass-through estimates following the 2018 U.S. tariff increases. Column (1) includes product-time and country-sector fixed effects. Column (2) includes product-time, country-time, and buyer-time fixed effects. The overall pass-through elasticity is computed as $1+$ the estimated coefficient on $\Delta\ln(1+\tau)$, and decomposed into contributions from the cost channel ($\Lambda_{ij}$) and the markup channel ($\Gamma_{ij}$). The counterfactual “Cost channel only” row shows the predicted pass-through when markup elasticities are set to zero. Panel B reports the relative contribution of each channel to the cross-sectional variance of $\Lambda_{ij}+\Gamma_{ij}$, the total elasticity governing pass-through. These shares sum to one and are derived from a variance decomposition. Source: FSRDC Project Number 2109 (CBDRB-FY25-P2109-R12520).
@@
Panel A of Table 7 reports aggregate pass-through predicted by the model, alongside a counterfactual that isolates the cost channel by shutting down markup adjustments. The model predicts aggregate pass-through around 73% under the baseline specification (Column (1)) and 67% under the specification with product–time, country–time, and buyer–time fixed effects (Column (2)), somewhat below the unweighted average of 78% reported in Column (2) of the top panel of Table 6, reflecting the greater weight placed on large, high-share buyers, which exhibit more incomplete pass-through.
Marginal cost adjustments account for most of the aggregate response. Shutting down markup elasticities ($\Gamma_{ij}=0$) yields pass-through rates between 71.5% and 64.8%, very close to the aggregate estimates. This indicates that the cost channel accounts for most of the aggregate response. Markup adjustments operate in the opposite direction, exerting only a small offsetting effect and therefore raising pass-through only slightly relative to the cost-only benchmark.
Panel B provides a complementary decomposition by analyzing the variance of $\Lambda_{ij}+\Gamma_{ij}$, which governs pass-through through $1/(1+\Lambda_{ij}+\Gamma_{ij})$. By construction, the contributions of cost and markup elasticities sum to one. Consistent with Panel A, nearly all cross-sectional variation in aggregate pass-through is accounted for by the cost elasticity $\Lambda_{ij}$, with markup elasticities playing only a minor role.
5.1 Comparison with Existing Pass-Through Estimates
Several influential studies document near-complete pass-through of the 2018 U.S. tariffs at the product level (e.g., fajgelbaum2020return; amiti2019impact; amiti2020s). Within our baseline sample of repeated firm-to-firm relationships, the implied average (unweighted) pass-through is bounded between 81 and 87 percent, substantially incomplete and closely matched by the model (see Appendix Tables 6 and D.7). [^32] In a specification comparable to those in the literature, Column (1) of Table 3 yields a duty-exclusive pass-through elasticity of 0.13-0.13 (s.e. 0.09); a more stringent set of fixed effects yields 0.19-0.19 (Appendix Table D.2).
Two features distinguish our setting from product-level studies. First, our baseline imposes sample restrictions: we focus on repeated, arm’s-length links trading intermediate and capital goods, with suppliers serving at least two buyers. Second, we measure prices at the buyer-supplier–product level and at annual frequency, whereas product-level studies aggregate across relationships and use monthly data. To assess how each contributes to the gap with the literature, we evaluate them in turn.
Sample restrictions.
To isolate the role of sample composition, we progressively relax our baseline restrictions while keeping the unit of observation at the relationship level. Appendix Table D.11 shows the results. Including consumption goods raises pass-through from 82% to 88%; allowing single-buyer suppliers raises it further to about 92%; and the broadest specification, which also restores related-party trade, energy goods, and extreme price changes, reaches 95%. Each restriction contributes, but none on its own closes the gap with the literature: within the sample of repeated supplier–buyer links, pass-through remains incomplete regardless of which other restrictions are relaxed.
Aggregation.
To isolate the role of aggregation, we aggregate our baseline firm-to-firm prices to the product–source-country level and use monthly rather than annual data, following the literature. Appendix D.5 provides more details on the data construction, and Table D.12 reports the results. Within our baseline sample of repeated links, pass-through remains incomplete even at the product level, confirming that aggregation alone does not close the gap. Similarly, pass-through remains incomplete even once all sample restrictions are lifted at the product level (Columns (1)–(4)). The remaining difference is accounted for by the final step: once we add non-consecutive relationships (Column (5)), that is, transactions not observed in both years, pass-through reaches completeness, matching the literature.
The near-complete pass-through in product-level studies is thus driven primarily by the inclusion of non-consecutive, one-off relationships. Within the population of ongoing buyer-supplier links, pass-through is systematically incomplete. Understanding aggregate pass-through therefore requires accounting for the extensive margin of trade, that is, how the set of active trading relationships responds to tariffs, beyond pricing within existing relationships.